OpenStudy (anonymous):

Can someone please help me? Choose the equation of the line passing through the point (2, 6) and parallel to y = -3x - 4. Options: y=3x-12 y=-3x y=-3x+12 y=3x

OpenStudy (anonymous):

@Gabylovesyou @ikram002p @ganeshie8 @IMHERETOHELP

OpenStudy (anonymous):

Ok, so we need to find the slope of y=-3x-4

OpenStudy (anonymous):

do u know slope formula?

OpenStudy (anonymous):

m=(y2-y1/x2-x1)

OpenStudy (anonymous):

but....

OpenStudy (anonymous):

we already know the slope

OpenStudy (anonymous):

m=-3

OpenStudy (anonymous):

So we can eliminate 2 answer choices right now, do u know which ones those r?

OpenStudy (anonymous):

A and D?

OpenStudy (anonymous):

yup!

OpenStudy (anonymous):

So.....

OpenStudy (anonymous):

substitute the value of x and see whcih one gives the the correct y output the will satisfy the given points

OpenStudy (anonymous):

that will be the answer!!!!!

OpenStudy (anonymous):

Can you help me with that because I am really confused on how to figure that out

OpenStudy (anonymous):

ok, so we have narrowed it down to B and C right?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

so the points that they gave us are (2,6)

OpenStudy (anonymous):

one sec

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

\[(2,6)---.(x=2, y=6)---> Answer B. y=-3x, y=-3(2), y=-6\]

OpenStudy (anonymous):

oh my mistake

OpenStudy (anonymous):

the answer is C but thats how u solve it lol

OpenStudy (anonymous):

ohhhh okay thank you so much. you showing me the steps has helped me know how you got it :)

OpenStudy (anonymous):

Do you have time to help me with one more?

OpenStudy (anonymous):

sure

OpenStudy (anonymous):

The equation below represents Function A and the graph represents Function B: Function A f(x) = x - 9 Function B

OpenStudy (anonymous):

which equation best compares the slopes of the two functions?

OpenStudy (anonymous):

Options: Slope of function B = 2 x Slope of Function A Slope of function A = Slope of function B Slope of function A = 2 x Slope of function B Slope of function B = - Slope of function A

OpenStudy (anonymous):

This is very easy, lol, what is the slope of Function A?

OpenStudy (anonymous):

Is it what it tells you right in the beginning of the question?

OpenStudy (anonymous):

The x=-9?

OpenStudy (anonymous):

=x-9 I mean

OpenStudy (anonymous):

ok, any time u just have a function that is like y=x+1, or y=x+2, the slope is 1 because there is only one x, if it is y=3/2x+1, then the slope is 3/2, so the slope of Function A is 1 because it is x-9, then as u read on in the question, it asks choice compares to the slopes of Function A and B, which means that function A and B have the same slope, which is 1. So B would be the best answer because the Slope of Function A is equal to the slope of Function B.

OpenStudy (anonymous):

wait

OpenStudy (anonymous):

let me check something

OpenStudy (anonymous):

Alright

OpenStudy (anonymous):

ok made another mistake lol

OpenStudy (anonymous):

ok so in response to your first question,, always remember that slope is always the number in front of x, if there is no number, then the slope is 1

OpenStudy (anonymous):

the slope of Function B is 2

OpenStudy (anonymous):

I used slope formula

OpenStudy (anonymous):

m=(y2-y1/x2-x1)

OpenStudy (anonymous):

m is slope

OpenStudy (anonymous):

I picked 2 x points and 2 y points

OpenStudy (anonymous):

m=(-3-1)/(-1-1)=2

OpenStudy (anonymous):

so the correct answer is C

OpenStudy (anonymous):

oh okay haha. Thanks for the help :)

OpenStudy (anonymous):

np